OpenStudy (anonymous):

find the first derivative with respect to x of y= 3x/8 + (sinxcosx)/8 +(cos^3 xsinx)/4 a.sin^4 x b.-sin^4 x c. cos^4 x d. -cos^4 x

OpenStudy (anonymous):

y= 3x/8 + (sinxcosx)/8 +(cos^3 xsinx)/4 y' = 3/8 + 1/8 * (cos^2(x) - sin^2(x)) + 1/4 * (-3cos^2(x)*sin(x) + cos^4(x) ) = = 3/8 + 1/8 * (cos^2(x) - sin^2(x)) + 1/4 * (-3cos^2(x)*sin(x) + cos^4(x) ) = = 3/8 + 1/8 * (1-2*sin^2(x)) + 1/4 * (-3(1-sin^2(x))*sin(x) + cos^4(x) ) = = 3/8 + 1/8 * (1-2*sin^2(x)) + 1/4 * (-3(1-sin^2(x))*sin(x) + (1-sin^2(x))^2 ) = = 3/8 + 1/8 * (1-2*sin^2(x)) + 1/4 * (-3(1-sin^2(x))*sin(x) + (1-sin^2(x))^2 ) = = 3/8 + 1/8 * (1-2*sin^2(x)) + 1/4 * (-3(1-sin^2(x))*sin(x) + 1-2sin^2(x)+sin^4(x) ) = = 3/8 + 1/8 + 1/4 - 3/4*sin(x) + (1/4-2/4)*sin^2(x) + 3/4*sin^3(x) + 1/4*sin^4(x) = .= 6/8 - 3/4*sin(x) -1/4*sin^2(x) + 3/4*sin^3(x) + 1/4*sin^4(x) = = 6/8 - 3/4*sin(x)(1-sin^2(x)) -1/4*sin^2(x) (1-sin^2(x)) = = 6/8 - 1/4*(1-sin^2(x))sin(x) (3 + sin(x))

OpenStudy (anonymous):

then ..

OpenStudy (anonymous):

i vote for none of the above

OpenStudy (anonymous):

ok this is the longest part of my h.w i will put E.